# Modeling 3D Infant Kinetics Using Adaptive Graph Convolutional Networks

### v1

@@ Abstract
Reliable methods for the neurodevelopmental assessment of infants are essential for early detection of medical issues that may need prompt interventions. Spontaneous motor activity, or ‘kinetics’, is shown to provide a powerful surrogate measure of upcoming neurodevelopment. However, its assessment is by and large qualitative and subjective, focusing on visually identified, age-specific gestures. Here, we follow an alternative approach, predicting infants’ neurodevelopmental maturation based on data-driven evaluation of individual motor patterns. We utilize 3D video recordings of infants processed with pose-estimation to extract spatio-temporal series of anatomical landmarks, and apply adaptive graph convolutional networks to predict the actual age. We show that our data-driven approach achieves improvement over traditional machine learning baselines based on manually engineered features.
@@ 3 Method
Given a sequence of infant skeleton recordings, we formulate the machine learning problem as the task to predict the age of the infant. Each recording is represented as a spatio-temporal graph, where nodes correspond to skeletal joints, and edges represent the physical connections between these joints. Let $\mathbf{X}\in\mathbb{R}^{C\times T\times V}$ be the input tensor, where $C=3$ represents spatial coordinates as node features, $T=600$ is the temporal dimension, and $V=18$, is the number of vertices, or joints in the infant graphs.
@@
The complete network architecture is shown in Fig. 2. Both the spatial part GCN (including STC for AAGCN), and the TCN are succeeded by a batch normalization (BatchNorm) [11] layer and a rectified linear unit (ReLU) [16] activation layer. To improve training stability and gradient flow, each block includes a residual connection [9]. Additionally, a dropout [24] of 0.5 is added between the GCN and TCN to mitigate overfitting. The network begins with a data BatchNorm layer to normalize the input data. The total number of STGCN or AAGCN blocks is set to 4. The output channel dimensions of these blocks are set at 32, 32, 64, and 64, respectively. Average pooling is applied on the output of the stack of spatio-temporal convolution blocks, which is further entered into a two-layer feedforward network with hidden dimensions 32 and 1, and a ReLU activation in-between.
@@
The training objective is to minimize discrepancy between model predictions and the actual age of the infants. Mean squared error (MSE), $\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}(\mathrm{a}_{i}-\widehat{\mathrm{a}}_{i})^{2}$, is used as the loss function. Here, $\mathrm{a}_{i}$ represents the actual age of the infant, $\widehat{\mathrm{a}}_{i}$ is the model’s prediction, and $N$ is the number of samples in each batch. The MoMo optimizer [22] with a learning rate of 0.01 was used to minimize the loss function. We employ $k$-fold cross validation with $k=10$ to get a robust assessment of model performance. In each fold, the models are trained for 20 epochs using a batch size of 32, where the best validation checkpoint is stored for each. The models are implemented in the PyTorch Lightning framework [7]. An Nvidia V100 GPU is used for training.
@@ 4 Results
[…]
@@
If we restrict model predictions to infants 10 to 20 weeks, a common age range for GMA studies, regression errors are clearly improved. AAGCN is still the best with RMSE = 11.12 ± 2.92 days. Overall, the prediction of age through the full cohort is challenged most notably by the youngest infants, who are known to express developmentally and qualitatively different selection of spontaneous movement patterns as compared to the typical GMA’s ‘fidgety age’ around three months of age [5].
@@
The learned graphs in each layer of the AAGCN model as it is adapted to an individual infant is shown in Fig. 3. The subject shown here is a normally developing infant, corrected postnatal age 135 days, that the model predicts as 144 days. Model inference takes 71 ± 12 ms on a CPU. Several new connections are formed in the skeletal graph changing the overall topology significantly. The strongest connections remain as edges from the original input graph, complemented by weak connections between joints further apart in the body. We can see that each layer in the model learn to focus on different body parts. Both the first and second layer emphasize on the lower legs and arms (2nd layer much more on the right arm) for the age assessment. The 3rd and 4th layers put more relative weight on the upper legs that the first two layers. All layers retain most of the bidirectional edges between hands and feet in the top $|\mathcal{E}|$ connections, showing that the coordinated movement of these joints is useful. Most focus overall is put on lower limb movement for determining the age here.
@@ 5 Conclusion
In this study, we introduced a data-driven approach utilizing pose-estimated infant recordings to predict age, highlighting the potential advantages of graph neural networks and the efficacy of three-dimensional poses for infant developmental screenings. We show that data-driven methods can extract the relevant features and outperforms machine learning based on limb movement parameters. Expanding the 3D screenings of infant general movements to an even greater scale would be immensely useful for future endeavours as deep learning is known to benefit from more data to enhance performance.

### v2

@@ Abstract
Reliable methods for the neurodevelopmental assessment of infants are essential for early detection of medical issues that may need prompt interventions. Spontaneous motor activity, or ‘kinetics’, is shown to provide a powerful surrogate measure of upcoming neurodevelopment. However, its assessment is by and large qualitative and subjective, focusing on visually identified, age-specific gestures. Here, we follow an alternative approach, predicting infants’ neurodevelopmental maturation based on data-driven evaluation of individual motor patterns. We utilize 3D video recordings of infants processed with pose-estimation to extract spatio-temporal series of anatomical landmarks, and apply adaptive graph convolutional networks to predict the actual age. We show that our data-driven approach achieves improvement over traditional machine learning baselines based on manually engineered features. The source code is available at https://github.com/deinal/infant-aagcn.
@@ 3 Method
Given a sequence of infant skeleton recordings, we formulate the machine learning problem as the task to predict the age of the infant. Each recording is represented as a spatio-temporal graph, where nodes correspond to skeletal joints, and edges represent the physical connections between these joints. Let $\mathbf{X}\in\mathbb{R}^{C\times T\times V}$ be the input tensor, where $C=3$ represents spatial coordinates as node features, $T=600$ is the temporal dimension, and $V=18$ is the number of vertices, or joints, in the infant graphs.
@@
The complete network architecture is shown in Fig. 2. Both the spatial part GCN (including STC for AAGCN), and the TCN are succeeded by a batch normalization (BatchNorm) [11] layer and a rectified linear unit (ReLU) [16] activation layer. To improve training stability and gradient flow, each block includes a residual connection [9]. Additionally, a dropout [24] of 0.5 is added between the GCN and TCN to mitigate overfitting. The network begins with a data BatchNorm layer to normalize the input data. The total number of STGCN or AAGCN blocks is set to 4. The output channel dimensions of these blocks are set to 32, 32, 64, and 64, respectively. Average pooling is applied on the output of the stack of spatio-temporal convolution blocks, which is further entered into a two-layer feedforward network with hidden dimensions 32 and 1, and a ReLU activation in-between.
@@
The training objective is to minimize the discrepancy between model predictions and the actual age of the infants. Mean squared error (MSE), $\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}(\mathrm{a}_{i}-\widehat{\mathrm{a}}_{i})^{2}$, is used as the loss function. Here, $\mathrm{a}_{i}$ represents the actual age of the infant, $\widehat{\mathrm{a}}_{i}$ is the model’s prediction, and $N$ is the number of samples in each batch. The MoMo optimizer [22] with a learning rate of 0.01 was used to minimize the loss function. We employ $k$-fold cross validation (split according to subject) with $k=10$ to get a robust assessment of model performance. In each fold, the models are trained for 20 epochs using a batch size of 32, where the best validation checkpoint is stored for each. The models are implemented in the PyTorch Lightning framework [7]. An Nvidia V100 GPU is used for training.
@@ 4 Results
The results of the 10-fold cross-validation are shown in Table 1. We report the average validation RMSE, MAE, MAPE, and R2 regression metrics across the folds, along with their standard deviations. The results indicate that the ensemble-based methods outperform the linear model in all metrics, with XGB being the best. The ensemble models also demonstrate a higher variance across folds compared to LR, suggesting a more sensitive response to specific folds.
@@
If we restrict model predictions to infants 10 to 20 weeks, a common age range for GMA studies, regression errors are clearly improved. AAGCN is still the most performant with RMSE = 11.12 ± 2.92 days. Overall, the prediction of age through the full cohort is challenged most notably by the youngest infants, who are known to express developmentally and qualitatively different selection of spontaneous movement patterns as compared to the typical GMA’s ‘fidgety age’ around three months of age [5].
@@
The learned graphs in each layer of the AAGCN model as it is adapted to an individual infant is shown in Fig. 3. The subject shown here is a normally developing infant, corrected postnatal age 135 days, that the model predicts as 144 days. Model inference takes 71 ± 12 ms on a CPU. All layers retain most of the bidirectional edges between hands and feet in the top $|\mathcal{E}|$ connections, showing that the coordinated movement of these joints is useful. This observation aligns with common clinical experiences where normal motor development is characterized by increasing coordination of movements across all limbs. Importantly, normal development requires two levels of coordination: First, the upper limbs and lower limbs need to be mutually correlated, such as hands touching each other and crossing the midline. Second, there is a need for coordination across levels, between upper and lower limbs, to prepare for full-body movements like crawling or walking. This layered coordination supports the hypothesis that coordinated limb movements are key indicators of neurodevelopmental maturity.
@@ 5 Conclusion
In this study, we introduced a data-driven approach utilizing pose-estimated infant recordings to predict age, highlighting the potential advantages of graph neural networks and the increased efficacy of 3D poses compared to 2D poses for infant developmental screenings. We show that data-driven methods can extract the relevant features and outperforms machine learning based on limb movement parameters. Expanding the 3D screenings of infant general movements to an even greater scale would be immensely useful for future endeavours as deep learning is known to benefit from more data to enhance performance.